A jar contains nickles and pennies. If there are 56 coins in all and the total value is $1.52, How many nickles are in the jar?

Please Help?

Found 2 solutions by ikleyn, rfer:
Answer by ikleyn(52810)
Answer by rfer(16322)

n+p=56
p=56-n
-------------
0.01p+0.05n=1.52
0.01(56-n)+0.05n=1.52
0.56-0.01n+0.05n=1.52
0.04n=1.52-0.56
n=0.96/0.04
n=24 nickels
56-24=32 pennies